OpenStudy (anonymous):

A cylinder has a radius of 14 m and a height of 6 m. What is the exact volume of the cylinder? A.1176π m3 B.504π m3 C.168π m3 D.84π m3

OpenStudy (anonymous):

volume = π · r2 · h volume = π · 14^2 · 6 volume = π · 196 · 6 volume = π · 1176 volume = 1176π m3 answer: A